OpenStudy (anonymous):

John has two rectangular mirrors that are similar. The length of the smaller mirror is four inches and the width is six inches. The width of the bigger mirror is two feet What is the length of the bigger mirror? Explain in words how you found your answer write why you took the steps you did to solve the problem.

OpenStudy (anonymous):

Let X2 be smaller mirror. X1 Larger mirror 1ft = 12inches X = LWH X2 = (4)(6)(h) X1 = (L)(24)(H) (4)(6)(H) = (L)(24)(H)

OpenStudy (anonymous):

L = 1

OpenStudy (anonymous):

i belivie u should times by two for the length of the small one to find out the length for the bigger one

OpenStudy (anonymous):

2ft is 24 in. *2and4 equal 8

Directrix (directrix):

Similar polygons have corresponding sides in proportion. Changing two feet to 24 inches gives 6 inches to 24 inches as the scale factor in common units of measurement. 6/24 = 1/4. Let w be the width of the larger rectangle. 6 is to 24 as 4 is to w. 6/24 = 4/w 1/4 = 4/w w = 16 inches or 16/12 = 1 and 1/3 feet.